Circular No. 8372

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004db IN NGC 7377 T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report that a spectrum (range 340–730 nm) of SN 2004db (“Possible Supernova in NGC 7377” announced on IAUC 8367), obtained by J. Huchra on July 10.46 UT with the F. L. Whipple Observatory 1.5m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova several weeks past maximum.

SUPERNOVA 2004dc IN IC 1504 Further to IAUC 8361, S. Park and W. Li report the LOSS discovery

of an apparent supernova (mag about 19.5) on KAIT images taken on June 29.46 and 30.40 UT. SN 2004dc is located at α = 23^{h}41^{m}18^{s}.47, δ = +4^{o}01′06′′.3 (equinox 2000.0), which is 14′′.8 west and 3′′.5 north of the nucleus of IC 1504. A KAIT image taken on 2003 July 1.45 showed nothing at this position (limiting mag about 20.0).

R. J. Foley, F. J. D. Serduke, W. Li, and A. V. Filippenko, University of California, Berkeley, report that inspection of CCD spectra (range 330– 1000 nm), obtained on July 11 UT with the Shane 3-m telescope at Lick Observatory, shows that SN 2004dc is of type Ic, probably several weeks past maximum brightness.

THE EDGAR WILSON AWARD The Smithsonian Astrophysical Observatory announces that the 2004

Edgar Wilson Award for the discovery of comets (cf. IAUC 6936, 7926, 8162) is being divided among the following two individuals: Vello Tabur, Wanniassa, A.C.T., for C/2003 T3; and William A. Bradfield, Yankalilla, S. Australia, for C/2004 F4.

V1186 SCORPII CCD magnitudes of V1186 Sco (cf. IAUC 8369, 8371) by B. E. Schaefer,

Louisiana State University, obtained with the 1.0-m SMARTS telescope at Cerro Tololo: July 6.97 UT, U = 11.37, B = 11.50, V = 10.79, R = 10.26, I = 9.73; July 7.22, 11.30, 11.40, 10.67, 10.15, 9.64; July 8.18, 10.95, 11.11, 10.36, 9.82, 9.33; July 9.23, 10.23, 10.37, 9.69, 9.19, 8.65; July 10.24, 10.79, 10.54, 9.75, 9.23, 8.71; July 11.20, 11.46, 10.92, 9.85, 9.23, 8.61.

2004 July 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green